Status of Diffractometers

Status of Diffractometers:

   Venture: OK

  D8 Duo: OK 


24 May 24:

The D8 Duo's copper shutter (micro-source) wouldn't open - Ariel from Bruker replaced the O-rings with the X-ray optics which solved the vacuum leak problem. The system was re-aligned and beams optimized and the instrument is now back in operation.

5 March 24:

The D8 Duo's copper shutter (micro-source) won't open - quick diagnosis points to a vacuum leak in the X-ray optics.  Waiting for further information on the repair from Bruker and will update with a timeline when one is known.

27 June 23:

The D8 Duo's detector warmed up overnight.  Some quick troubleshooting (thanks to Ariel) and we found that a fuse inside the APEXII power supply (for the Peltier cooling block had blown.  A quick replacement and all seems well.

5 May 23:

The chiller pump inside the Lytron has been replaced, along with a cracked fittingNo leaks overnight so it seems to be good to go.

3 May 23:

Looks like the chiller pump inside the Lytron has sprung a leak.  We have a spare on the shelf (hopefully the seals are still good) so Dirk will work on replacing it.

23 Feb 23:

Ylid dataset looks good so the Platform is back in operation.

16 Feb 23:

Apparently we are only allowed to have one operational system at a time.  Platform stopped during a data collection and lost its connection to the GGCS.  The control program GC is giving an error:  

Error in GC_PARMS()

Scaler already disabled

We checked all of the boards inside the GGCS, batteries, etc.  Ended up swapping in a different scaler board from a donor system and communication has been restored.  Tried to home the axes - phi and 2theta are good, but trying to move omega gave a nasty sound.  Looks like the omega drive motor is shot.  We are working at replacing with a motor from our donor system, and also giving the ring gear for omega a good cleaning.  Big thanks to Farhat from our electronics shop and Dirk from our machine shop for taking the lead on the repair.

26 Jan 23:

Update on the D8.  Ylid datasets look great with both the Mo and Cu beams. Tested a real-world sample and everything looks good.  Back to normal operation.

24 Jan 23:

Update on the D8.  Ariel from Bruker is on-site to align the consolidated beam path.  Fingers crossed that we are fully operational by the end of the week.

5 Oct 22:

Update on the D8.  We received and installed a new PCB for the Cu high speed shutter - we now have an operational shutter.  The downside is that the consolidated beam path is now out of alignment; awaiting further instructions from Bruker.

12 Aug 22:

Huge thanks to Farhat from our electronics shop!  We rebuilt the water flow meter but still couldn't get the generator back on.  A little more troubleshooting and we found that the solenoid-controlled water valves (bypass and cooling circuit) were not operating properly.  Farhat got them cleaned up and we seem to be back in business.  It's going to take at least two weeks to work through the backlog, so please be patient.

8 Aug 22:

Don't go on vacation!  Upon returning from a week off, I discovered that the platform generator had tripped off.  Looks like an issue with the cooling water flow sensor on the generator.  Parts have been ordered from Bruker.

18 Jul 22:

After swapping three different controller boards to no avail, it looks like the fault with the microsource shutter - either the shutter motor and/or controller PCB.  We are going to pull the assembly out for further testing of the motor in our electronics shop before ordering the replacement bits.  As an added bonus, the Haskris water pump (for cooling the Mo generator and tube) has seized so now we are waiting on parts from Haskris.

25 Feb 22:

A couple different issues with the D8.  First, the Cu microsource shutter is not opening.  Second, the Mo side of the diffractometer seems to also be experiencing issues; both problems point towards a failing board in the D8 controller or the enclosure.  Parts have been ordered.

06 Oct 21:

The long awaited update. We received an exchange generator from Bruker for the copper microsource. After a couple minor problems, we've got the instrument back up and running. There is a bit of queue now so there will be some delays in working through the backlog. Thanks for your patience.


05 Jul 21:

With the guidance of Kerry and Todd from Bruker, I was able to align the D8. The cause of the collision was a broken cable that connects the crash plates to the goniometer (in theory the crash plates should have detected that a collision was imminent and halted the goniometer - that didn't happen because the cable was broken). A big thanks to Farhat from our electronics shop for repairing the cable. We've run into another problem, which is the Cu generator is overshooting its setpoint when ramping to full power and thus the generator is not stable. Bruker has contacted Incoatec for more information.


18 Jun 21:

COLLISION - it appears that the collision plates on the D8 failed to do their one and only job - the chi arm appears to be out of alignment following a collision with the detector. Working with Bruker to find a solution.


13 May 21:

Platform is back in operation!


12 May 21:

TESTING - Problem solved! Farhat replaced the phi driver board located inside the GGCS from our donor system. Also discovered that the detector enable cable that connects to the GGCS was broken right at the connector. Farhat fixed this too and now the diffractometer seems to be okay. Will run a test dataset overnight.


11 Feb 21:

The communication problem has been resolved but now a new issue - the phi axis does not seem to move at all (omega and 2theta are fine). It's likely the phi cable, the phi interconnect board (on the underside of the goniometer), the phi driver board inside the GGCS, or the phi motor itself.


04 Feb 21:

There is a communication problem betwee the PC controlling the instrument and the GGCS. Farhat (from electronics) is replacing the ancient batteries surface mounted to the motherboard of the GGCS. Fingers crossed that this solves the problem.


30 Jun 20:

Leak seems to be repaired (thanks to Dirk from our machine shop!). Detector is back at operating temperature. Will be keeping a close eye for any new problems.


29 Jun 20:

Discovered a coolant leak in the Lytron chiller (that is responsible for cooling the detector). Need to order in some new fittings and hoses.


27 May 20:

Leak has been fixed and the diffractometer is operational.


19 May 20:

Small water leak on the building supply line that connects to the Haskris for the Platform diffractometer. Work order is in with the university plumbers so just awaiting our turn in the queue.


24 Mar 20:

As of 5pm, Tues, March 24, 2020, the department will be closed due to COVID-19. Both diffractometers will be powered off and unplugged.


13 Feb 20:

New Mo tube installed and we are back in service.


24 Jan 20:

Looks like the Mo tube on Platform has failed after 5 years of service. Time to order a new one!


09 Oct 19:

The Kryoflex seems to be functioning ok, not 100%, but has successfully run overnight without condensation building up externally nor icing up internally. We have resumed collecting data on this instrument now but caution that there may still be issues down the road.


20 Sep 19:

Quick update on the D8 - the power supply inside the IuS generator was replaced and we also discovered that the two fuses at the power connector had also blown. Generator is now back operational. Still troubleshooting the Kryoflex.


3 Sep 19:

A couple problems that we are dealing with: first, there is a gas flow issue with the Kryoflex that is leading to icing at the top of the gas out port of the heat exchanger. We are troubleshooting a rebuilt pump while we await the arrival of the new pump. Second, it seems as though the power supply inside the IuS generator has failed. A new one is on order.


21 Jun 19:

Turns out the motor for the omega drive was not repairable; we salvaged a motor and gear drive from a donor system and are now back operational. Once again, many thanks to our machine shop and electronics shop for their help.


14 Jun 19:

And now an issue with the omega drive motor of the Platform, as it appears that the bearings internal to the motor are failing. Our machine shop is handling the repair.


24 May 19:

We received a donated BVT3300 Kryoflex controller from Bill Brennessel (Univ Rochester) for the D8 and the instrument is once again operation. Many thanks to Bill and the Bruker Users group for their assistance.


01 Apr 19:

An issue with the Kryoflex temperature controller on the D8 system following the weekend power shutdown. All other components have come back online properly. Consulting with Bruker and will update this page as information becomes available.


02 Nov 18:

The cooling fan in the base of the platform goniometer needed to be replaced. Thanks to our electronics shop and machine shop for their assistance. Currently running a test crystal over the weekend.


12 Oct 18:

The D8 is back up and running after a full alignment of both sources by Bruker. Huge thanks to the department for its support of our facility.


05 Oct 18:

The Bruker service engineer is expected to arrive on Tues, Oct 9, to re-align the D8.


10 Aug 18:

Looks like our copper beam is out of alignment following a collision. Will again be consulting with Bruker.


21 Jul 18:

After a couple of Webex sessions with Bruker we narrowed the problem down to either the phi interconnect board or the phi cable. We pulled a cable from a donor system and installed on the D8 - seems to be working okay now.


27 Jun 18:

There is an issue with the phi-drive of the D8 diffractometer. Troubleshooting with Bruker.


6 Feb 17:

The Cu microsource is back up and running. Thanks to Bruker for walking us through the replacement and alignment process. As always, major thanks to our department for its continued support of our facility.


20 Jan 17:

After consultation with Bruker we've determined that the Cu microsource X-ray tube has burned out. We've ordered a replacement and hope it should get to us by early March.


29 Nov 16:

The Cu microsource on the D8 Duo diffractometer has failed. Consulting with Bruker to see what exactly is the problem. In the meantime, we can at least use the Mo sealed tube source on this machine, but it means that we won't be as able to get data from small weakly-diffracting crystals. But we will try whatever gets submitted.


24 Aug 16:

The D8 Duo's low-temperature system is back in working order thanks to Dirk in our Machine Shop.


22 Aug 16:

And now it's the low-temperature system on the D8 Duo machine. Replaced a gas pump and the gas couplings aren't fitting so well. At least there's a local supplier for these. Hope to be up before the end of the week.


12 Aug 16:

The D8 Duo system is back in normal operation.


11 Aug 16:

The pump repair kit arrived and we seem to be back in business. As usual, we'll do an overnight test to see if the system is stable.


9 Aug 16:

STILL WAITING on that parts replacement kit for that vacuum pump. In the meantime we've ordered a kit from another supplier who should be able to get it to us more quickly. At least we'll have a replacement set on hand for the next time it wears out.


13 July 16:

Problems with an air pump that evacuates the beam path through the focusing mirrors for the Cu microsource on the D8 instrument. We've ordered a servicing/rebuild kit and hope it gets here in a few days.


9 Dec 15:

No problems with the D8 Duo since we turned it back on over a week ago, so we'll say we're back to normal operation on this machine.


1 Dec 15:

We've installed a new high-voltage cable and replaced the burnt-out X-ray tube on the Mo-wavelength generator of the D8 Duo machine. Now running data collections on this instrument and keeping an eye on the stability of the generator.

18 Nov 15:

After much testing we've narrowed the problems with the D8 Duo's Mo generator down to either a faulty high-voltage cable or a malfunctioning transformer tank. The former is easily fixed (just buy a new cable!) but the latter means we'd have to ship the generator unit back to Bruker (Madison) for servicing. We've ordered a new cable along with a new Mo tube, which we need either way. Hope to have these for testing some time next week. Special thanks again to Al Chilton for his electronics wizardry.


20 Oct 15:

The good news is that the replacement generator unit for the Cu microsource is installed and is functional. The bad news is that the generator for the Mo sealed-tube source is acting up, probably due to a damaged high-voltage cable, which also damaged the tube in the process. We're doing further tests to determine whether other parts of the generator were also affected, and to figure out what replacement parts need to be ordered.


28 Sep 15:

We'll be sending the Cu microsource's generator unit back to Bruker in exchange for a working one. Unfortunately that means the D8 diffractometer will be totally out of service, since its safety interlocks are now tied to having two functioning generators. We do hope to have the new Cu generator control module installed on our system before the end of October.


26 Aug 15:

More problems with the D8 Duo machine. This time it's the Cu microsource generator, which seems to be unable to reach full operating power. Hoping the Bruker folks have some advice for us.


21 Aug 15:

Again our Electronics and Machine Shop wizards bail us out. The D8 diffractometer is back in operation.


14 Aug 15:

Problems with one of the axial positioning drives on the D8 machine. Hope to have this figured out in a few days.


21 July 15:

The PLATFORM diffractometer has passed its tests and is back in service. Many many thanks to Al Chilton (Electronics Shop) and Dirk Kelm (Machine Shop) for again getting us back in production.


20 July 15:

The PLATFORM machine is back in one piece and we're running a test dataset with a standard crystal. Hope to be back in normal operation within the next day or two.


14 July 15:

Not only a burnt-out fuse on the 2θ axis control board, but worn-out bearings on the driveshaft assembly for this axis. Al from the Electronics Shop and Dirk from our Machine Shop are on the case.


10 July 15:

More PLATFORM system woes, this time it's the goniometer's 2θ drive that's unresponsive. Burnt-out drive motor? Faulty logic board? We'll hand this one over to our elite electronics experts.


2 July 15:

The PLATFORM diffractometer is back up and running with the aid of the 21-year-old refrigerated water chiller that we used to use with the now-defunct rotating-anode X-ray generator.


29 June 15:

After a scheduled electrical shutdown in the department this weekend, the water recirculator pump (for cooling the X-ray tube) on the PLATFORM machine has decided to fail. Off to our machine shop for repairs.


2 Jan 15:

The PLATFORM diffractometer is back in normal operation.


22 Dec 14:

The new tube for the PLATFORM diffractometer arrives just before Xraymas. Now for the alignment process.


3 Dec 14:

A burnt out X-ray tube for the PLATFORM diffractometer. A replacement is on order and should arrive in a few days.


23 July 14:

A problem with the D8's generator control board. Another fix by Al from the department's electronics shop. Thanks once again to Bruker for their help.


18 July 14:

Problems with the Mo-wavelength generator on the D8 Duo diffractometer. Working with Bruker to troubleshoot.


30 May 13:

The D8 Duo diffractometer is back in regular service.


28 May 13:

The replacement Mo tube arrived this afternoon. Now to install it on the D8 Duo and go through the beam alignment steps. Should take a day or two.


17 May 13:

The molybdenum X-ray tube on the D8 Duo system has burnt out, which is to be expected after about five years of service. A replacement has been ordered, which should take a week or so to arrive. Unfortunately we can't use the copper microsource on this instrument since the safety interlocks are heavily dependent on the status of the original Mo sealed-tube generator.


2 Apr 13:

Another great job by our Electronics shop (thanks Al!). The PLATFORM system (now known as Old Sparky) is good to go.


27 Mar 13:

Some problems with the high-voltage power inputs on the PLATFORM system. Another job for our Electronics wizards. May need some rewiring of the power distribution module, but we hope that should only take a couple of days.


16 Jan 13:

Just a blown fuse on one of the axial-drive control boards (thanks again to Al from our Electronics shop). The PLATFORM diffractometer is back in operation.


15 Jan 13:

Trouble with one of the drive axes on the PLATFORM system. Hope to have it back in service by the end of the week.


12 July 12:

The PLATFORM system has been repaired and is back in production.

9 July 12:

After a building water shutdown on Saturday, problems with getting the generator on the PLATFORM system going. Sounds like it may be a blown capacitor on a main control board, so that's a job for our electronics folks.

10 Aug 11:

The (newly-redesignated) D8 Duo system is in normal operation.

28 July 11:

The PLATFORM system is back in normal operation.

The copper microfocus source has been installed on the D8 system. Right now we can't use the original molybdenum source on this machine due to interference between the beamstop and the low-temperature system's cooling nozzle. Bruker is working on a solution and we hope to have the D8 running in dual-wavelength mode later next week. Meanwhile we can collect datasets with the copper source, so after some test datasets with standard crystals we hope to start normal data collections next week.

27 July 11:

The Bruker service people are installing the copper microfocus source on the D8 system. After the alignment and calibration steps, we hope to have this one running by the end of the week.

Parts for the broken refrigerated chiller have arrived, so we hope to have this one hooked up to the PLATFORM system in the next day or so.


18 July 11:

In preparation for next week's addition of the copper microfocus source to the D8/APEX II system, we've taken the refrigerated chiller that was being used with the PLATFORM system and attached it to the D8's detector. Hoping to have the PLATFORM system back in service by the end of the week or early next.


17 July 11:

A series of problems with the D8/APEX II system, starting with the tube shutter, followed by the tube-cooling water recirculator, and now the refrigerated chiller that keeps the APEX II detector's CCD chip cold. Many thanks to Al Chilton of our Electronics Shop for his repair of the shutter, and Paul Crothers of our Machine Shop for working on the water recirculator. The refrigerated chiller is in the hands of our university's refrigeration systems specialist, and we're waiting on a small part which should get us running again in the next week or so.

1 June 11:

We will be upgrading our D8/APEX II system to dual-wavelength capability (Mo/Cu) with the addition of a copper microfocus source. Installation is scheduled for the last week of July, so we hope to have it in normal service at the beginning of August.

We are grateful to NSERC for funding of the equipment, and to the departmental users group for support of our application.


3 Feb 11:

The upgraded PLATFORM/APEX II system is in service.

28 Jan 11:

We will be replacing the old SMART 1000 detector on the PLATFORM system with a new APEX II detector. Installation is scheduled for the first week of February.

3 Sept 10:

Al from Electronics saves the day again. One replaced microswitch and we're back in production.

2 Sept 10:

Some problems with the D8/APEX II's shutter assembly. Al from our Electronics Shop is on the case.

4 May 10:

Back to normal operation.

3 May 10:

New gas pump received and installed. If it gets through the night with no problems we'll be back in business.


26 Apr 10:

Looks like the pump on the D8/APEX II's low-temperature (crystal cooling) system is dead. Replacement ordered; hope to be operational later this week or early next.


31 Aug 09:

The repairs to the D8/APEX II system seem to have held up over the weekend, so it's back in production.


28 Aug 09:

Repairs to the D8/APEX II system's generator have been completed. We're keeping a close eye on it and hope to be fully operational early next week.

Early results from the SMART 1000 detector looked promising after we had pumped down the mini vacuum chamber around its CCD chip. Unfortunately the chip again looks like it can't maintain its usual operating temperature (below -40 °C) so we may have to take it back out of service.

19 Aug 09:

The problem with the APEX II detector has been fixed, but now the X-ray generator won't fire up. We're working with Bruker and our in-house electronics people on this one, and in the meantime are also working on the old SMART 1000 detector, which again seems to have trouble keeping its CCD chip cold enough for data collection.


5 Aug 09:

A blown pump on the chiller that keeps the APEX II detector chip cold. Hoping for a replacement by some time next week.


22 June 09:

Some problems with the department's in-house nitrogen gas supply system, which is used by the low-temperature systems for the diffractometers. The supplier is on the case.

UPDATE: Problem solved. Back in action.


3 June 09:

The D8/APEX II system is back in normal operation.


1 June 09:

A new pump for the low-temperature system seems to have done the trick. We'll see how it goes overnight.


25 May 09:

Some (more) problems with the low-temperature system on the D8/APEX II system; back to the old PLATFORM/SMART 1000 machine for a while.


1 Apr 09:

We've restored the new D8/APEX II system to operational status.


27 Mar 09:

After a couple of different approaches, it looks like the low-temperature system is back in good health. We'll keep an eye on it over the weekend, and next week our Bruker service engineer will be here to check our work.


17 Mar 09:

We're working with Bruker to resolve the problem with the new machine's low-temperature system. Meanwhile we've reverted to data collections on the older system.


16 Mar 09:

Some issues with the low-temperature system on the (new) D8/APEX II system. We may have to shift back to the old PLATFORM/SMART 1000 machine while we work on it.


1 Aug 08:

The new D8/APEX II system is in normal operation.


29 July 08:

Looks like the low-temperature system is almost sorted out. We'll need to run a test dataset or two, and hope to be collecting datasets on the new machine by late this week/early next.


24 July 08:

The new diffractometer is in our lab and has been assembled. Unfortunately a key component of the low-temperature (crystal-cooling) system was damaged during shipment, and it may be later next week before we receive and install a replacement. The new software for diffractometer control and data processing is also totally different from what we've been using on our current machine, so it will take us some time to get familiar with it. Fortunately the old machine is still working.


12 July 08:

Update on the new diffractometer installation: it'll be the week of 21 July (not 14 July as previously mentioned).


7 July 08:

Shutter and water recirculator fixed. Back in business.


4 July 08:

Some problems with first the tube water recirculator, and now the X-ray shutter assembly. Hope to be back up in a couple of days.

If there's a silver lining, it's that we expect to receive our new diffractometer system by the end of next week. A Bruker service engineer will then complete the installation and testing of the new system, and we hope to be running samples on it before the end of the month.


5 June 08:

We were successful in our recent application to NSERC for funding for a new CCD diffractometer. We also thank the Faculty of Science at the University of Alberta for their generous partial support of the purchase of this instrument. The new machine has been ordered, and we hope to have it installed by late summer or early autumn.


24 Apr 08:

An anniversary of sorts, as we're marking one (trouble-free) year since our CCD detector went back into operation.


24 Apr 07:

The detector passed its tests, so it's back in production.


23 Apr 07:

Our detector has returned from Bruker and we've reinstalled it on our system. A couple of unexpected problems have occurred, so it may be a couple of days before we're back in normal operation.

UPDATE: Not as bad as we originally thought. Watch this space!

2 Mar 07:

We're back to normal (more or less).


1 Mar 07:

The loaner system has arrived and is on the diffractometer. Now for the standard tests - hope to start collecting datasets tomorrow.


26 Feb 07:

No detector yet - should be here by the end of this week.


14 Feb 07:

We've sent the CCD detector back to Bruker, but they've located a loaner unit for us. We should have it by the end of next week (around Friday 23 February) and should be back in production shortly thereafter.

1 Feb 07:

The CCD detector is again losing its ability to keep the CCD chip cooled to its normal operating temperature of -45 °C. It will probably have to be sent back to Bruker within the next week or so. We're trying to get through the samples we have in hand, but PLEASE don't send us any new ones until we find out whether we'll be able to get a detector on loan from Bruker (there are very few loaners available, and they're always in high demand).


25 Jan 07:

We're back in business.


23 Jan 07:

Got the shutter problem sorted out, but we had to move the tube quite a bit to get at it. Now for realignment - should only take the next day or so.


22 Jan 07:

A problem with the X-ray tube's shutter assembly. We'll keep you posted.


20 Nov 06:

We've received our detector back from Bruker after repairs. We'll be putting it back on our diffractometer and doing alignment and testing today, and running a dataset with the standard crystal overnight. We should be back in normal operation tomorrow.

Many thanks to the group at Bruker for arranging the loaner detector system for us.


24 July 06:

The loaner detector system worked well over the weekend, so we're back in normal operation.

We've put the number of samples in the waiting list on the lab home page to give our users an idea of how busy we are (or not). Past experience leads us to estimate that we can handle 12-14 samples per week, on average (including both successful data collections and unsuccessful sample trials).


20 July 06:

The 'loaner' CCD detector has arrived and has been installed on our diffractometer. We'll run a dataset on our standard test crystal, and with any luck we'll start data collections over the weekend.


17 July 06:

The 'loaner' system has passed its tests at Bruker (Madison) and has been shipped. We should have it in the lab by the middle of this week.


11 July 06:

Word from Bruker this morning is that the work on our 'loaner' system is proceeding, and that they should be finished refurbishing and checking it by the end of this week. If all goes well, we'll have it in our hands next week (week of 17 July).


29 June 06:

Bruker will taking back our detector to send back to Photometrics to see what went wrong. In the meantime they're assembling a detector system to send to us on loan. Because it needs to be configured and tested it'll probably get to us the week of 10 July.


24 June 06:

After three days' normal operation, the detector is unable to keep the CCD chip cool. Back to Bruker for advice.


20 June 06:

The CCD detector arrived back in our lab this morning. It's back on the diffractometer and is undergoing testing. Give us a day or two to give the 'all-up' signal.


9 June 06:

Still waiting on the CCD detector.

The newly-repaired anode seal on the rotating-anode generator has failed after little more than two weeks in service, so our 'backup' is also down.


29 May 06:

The CCD detector is back at Bruker after the repairs by Photometrics. Bruker now have to do a full recalibration on the detector, which will take the rest of this week and probably part of next. Optimistically we may get the detector back by the end of next week, but more realistically we should get it back in the lab some time during the week of 12 June.


23 May 06:

Some minor good news with the P4/RA. The repaired rotating-anode target assembly arrived back from Bruker today, so we've begun the (tedious) pump-down/power-up/align process. Hoping to have crystals on the machine by the end of the week


30 Mar 06:

More problems, but this time with the P4/RA. The rotating-anode vacuum seal has failed, and we've shipped the anode back to Bruker for a rebuild (it's not a user-serviceable part). Although this anode had only been in service for little over a month in 2000 (before we shifted the CCD detector to the sealed-tube generator), the vacuum seals tend to have a finite shelf life.

14 Mar 06:

Word from Bruker is that the repair price is quite reasonable. The detector is on its way to Photometrics Inc. for a rebuild. Estimated time: 2-3 months, due to the age of our detector, the complexity of the repairs, and their service queue.

We have been collecting data using the P4/RA, but it's slow going. Datasets typically take 2-3 days - sometimes longer if the generator goes down overnight, when we're not here to restart it.


1 Mar 06:

(No new news about the CCD detector, but in the meantime...)

Some encouraging signs as we try to get the old P4 diffractometer/rotating anode generator combo back in operation with the old-style serial (single reflection) detector. Many thanks to Al Chilton in the electronics shop for spotting and fixing three different problems with the diffractometer control interface. Now for the tedious alignment process.


27 Feb 06:

It looks like the cooling unit inside the vacuum chamber has failed, as Bruker were unable to get the CCD chip to cool below -25 °C (normal chip operating temperature is -45 °C). The camera will have to be sent back to the manufacturer (Photometrics Inc.) for repair or replacement of the Peltier cooling unit. Estimated time for repair: 8-12 weeks (combination of the job being complex and their service queue being lengthy). We also don't yet know how much the repairs would cost, and it's difficult for Bruker to give us a precise estimate since this doesn't happen often. Right now they do have a unit similar to ours at Photometrics for repairs. They hope to get it back within the next few days, and should then be able to give us an idea of the repair charges


21 Feb 06:

Bruker were able to re-establish the vacuum in the CCD chamber. They'll be keeping a close eye to see that it doesn't deteriorate (i.e. that the vacuum seal is still good). They then will commence testing the detector performance before okaying it to be shipped back to us.


13 Feb 06:

We've sent the CCD detector back to Bruker in Madison. The unit appears to have lost vacuum in the chamber containing the CCD chip and the cooling unit. This results in unusable data since the CCD chip isn't able to be cooled to its recommended operating temperature. The Bruker people will re-establish the vacuum (10-6 torr) and will be checking to see whether the seal has failed (and will repair if necessary). With any luck, we should be back in operation within two weeks.